# Hybrid Meta-learners for Estimating Heterogeneous Treatment Effects

### v1

@@ 4 Related Work
4.1 Performance Trade-off Between Indirect and Direct Learners
As discussed in Section 2, the regularization strategies employed by direct and indirect learners reflect different inductive biases, leading to distinct strengths and weaknesses. This divergence in performance has been consistently reported across the literature. In particular, prior work has shown that no single meta-learner consistently outperforms others across all settings [5]. Both theoretical and empirical studies demonstrate that their relative performance can vary significantly depending on the underlying DGPs, quantified by factors such as sparsity, smoothness, and the relative complexity of the POs and CATE functions [18, 10]. Acharki et al. [19] further compared meta-learners in the multi-treatment setting, theoretically revealing a bias–variance trade-off shaped by their estimation strategies. These performance trade-offs highlight the importance of developing a more robust approach, as the true DGP is often unknown in practice, and relying on either method can result in unreliable performance.
4.2 Regularization Strategies for Indirect Learners
Several other regularization strategies have been proposed to mitigate “regularization-induced confounding” in indirect learners. Weight regularization penalizes discrepancies between the weights of the outcome heads to encourage parameter sharing [11, 10]. Reparameterization implicitly introduces regularization by modeling $\mu_{1}(x)=\mu_{0}(x)+\tau(x)$, allowing the treatment effect $\tau(x)$ to be predicted directly by one head as an offset from $\mu_{0}(x)$ [10, 20]. Structural regularization was recently introduced with FlexTENet, a specialized neural architecture that enforces regularization through flexible feature sharing across layers [10].
@@ 5 Experiments
In this section, we evaluate the performance of H-learner through semi-synthetic experiments and established benchmarks. Section 5.1 outline the datasets and experimental setup. In Section 5.2, we highlight the performance trade-offs between indirect and direct learners using custom-designed semi-synthetic experiments and demonstrate the robustness of H-learner. In Section 5.3, we present results on widely used benchmark datasets, further highlighting the improved performance of H-learner over existing baselines.
@@ 5.1 Setup
Datasets. We evaluate performance using two widely adopted benchmark datasets: IHDP and ACIC 2016, both commonly used in the literature on heterogeneous treatment effect estimation [2, 10, 21, 18, 22]. For both datasets, we use a 63/27/10 train/validation/test split.
IHDP. The IHDP dataset is based on covariates from the Infant Health and Development Program, a randomized study assessing the effect of specialist home visits on cognitive outcomes for premature infants [14]. Confounding and treatment imbalance are introduced by removing a subset of treated units. The resulting dataset includes 747 observations with 25 covariates. Following prior work, we evaluate performance by averaging over 1,000 realizations from setting “A” implemented in the NPCI package [23].
ACIC 2016. The ACIC 2016 datasets are derived from the Collaborative Perinatal Project and were introduced as part of the Atlantic Causal Inference Competition [24]. This benchmark provides a comprehensive evaluation that varies in functional complexity, confounding, covariate overlap, and treatment effect heterogeneity. It includes 77 distinct data-generating processes, with 4,802 observations and 79 covariates. Following [22], we remove rows where the POs fall above the 99th percentile or below the 1st percentile to eliminate outliers and evaluate performance by averaging over 5 random realizations per setting.
Evaluations. We evaluate performance using Precision in Estimation of Heterogeneous Effect (PEHE) [14]. The PEHE loss is defined as $\epsilon_{\text{PEHE}}=\frac{1}{n}\mbox{$\sum_{i=1}^{n}$}(\widehat{\tau}(x_{i})-\tau(x_{i}))^{2}$, which corresponds to the mean squared error in estimating the CATE. We report performance on both in-sample (training and validation) and out-of-sample (test) data. Note that in-sample evaluation is non-trivial, as only the factual outcomes are observed during training, while the true CATE $\tau(x)$ is never observed.
@@ 5.2 Semi-Synthetic Experiments
Experiment setup. In all setups, we use covariates from the IHDP dataset ($n=747,d=25$) but design custom response surfaces as follows.
Let $S_{0}$ and $S_{1}$ denote the sets of randomly selected features for $\mu_{0}(x)$ and $\mu_{1}(x)$, and let $S=S_{0}\cup S_{1}$. We simulate the response surfaces and treatment assignment as:
$Y_{i}=2\,\mbox{$\sum$}_{\begin{subarray}{c}j\in S_{T_{i}}\end{subarray}}X_{j}+2\,\mbox{$\sum$}_{\begin{subarray}{c}j\in S_{T_{i}}\end{subarray}}X_{j}^{2}+\mbox{$\sum$}_{\begin{subarray}{c}j,k\in S_{T_{i}}\end{subarray}}X_{j}X_{k}+\varepsilon_{i},\quad\mathbb{P}(T_{i}=1\mid X_{i})=\sigma\left(\alpha\cdot\mbox{$\sum$}_{\begin{subarray}{c}j\in S\end{subarray}}\beta_{j}X_{j}\right)$
@@
In Setup A, we vary the relative complexity of the CATE and PO functions by changing the proportion of shared features between $S_{0}$ and $S_{1}$ in $\{10\%,30\%,50\%,70\%,90\%\}$, where a greater number of shared features implies a simpler CATE function. In Setup B, we vary group imbalance by adjusting the proportion of treated units in the dataset to $\{20\%,30\%,40\%,50\%\}$, where a higher percentage indicates more balanced treatment and control groups. In Setup C, we introduce varying levels of confounding in the treatment assignment function by adjusting $\alpha\in\{0,0.2,0.4,0.6,0.8\}$, where larger values of $\alpha$ correspond to stronger confounding.
In all setups, we fix the sizes of $S_{0}$ and $S_{1}$ to 10. To ensure that only one factor varies at a time, we set $\alpha=0$ in Setups A and B, and fix the proportion of shared features at 40% in Setups B and C. All simulations are averaged over 80 runs, with features in $S_{0}$ and $S_{1}$ randomly selected in each run.
Figure 4: PEHE results for three semi-synthetic scenarios. In all setups, we observe performance trade-offs between TARNet (indirect) and X-learner (direct), while the proposed H-learner consistently outperforms both counterparts.
Results. Figure 4 presents results from the three simulated scenarios. We show the performance of the H-learner using X-learner pseudo-outcomes in its first stage and compare it against its constituent components: the X-learner (direct learner) and TARNet (indirect learner). Results using alternative direct learners are provided in the supplementary materials.
Setup A demonstrates the performance trade-off through the relative complexity of the POs and the CATE. As the proportion of shared features increases, the PO functions become more similar, which leads to a simpler CATE surface. Correspondingly, we observe that the X-learner only outperforms TARNet when the proportion of shared features is high, as its constructed pseudo-outcomes target a simpler function, but underperforms when the shared feature proportion is low.
Setup B demonstrates the performance trade-off driven by the data imbalance between treated and control populations. As the proportion of treated units decreases, the imbalance in sample sizes between the treatment and control groups becomes more severe. This imbalance disproportionately affects indirect learners, as their separate outcome models are trained on increasingly unbalanced samples and are more susceptible to “regularization-induced confounding”. Empirically, TARNet performs better under balanced treatment assignment but underperforms when only 20% of units receive treatment.
@@
In contrast, across all three setups and varying parameters, the H-learner consistently outperforms both its direct and indirect counterparts by dynamically adjusting its inductive bias to better align with the underlying data, demonstrating robustness by unifying the strengths of both learners.
We further investigate the performance of H-learner across different $\lambda$ values. Figure 5 shows how the test PEHE varies with $\lambda$ in Setup A, across different levels of feature sharing: low (10%), medium (50%), and high (90%).
Results show that the optimal value of $\lambda$ consistently lies between the two extremes—$\lambda=0$ (corresponding to the indirect learner) and $\lambda=1$ (corresponding to the direct learner)—further highlighting that combining both regularizations yields better performance than relying on either alone. Across different levels of feature sharing, the H-learner with the best-performing $\lambda$ consistently improves upon the weaker learner by at least 10% and the stronger learner by at least 5%. Moreover, the trend in the optimal $\lambda$ reflects the inductive bias inherent in the data: as more features are shared, the underlying CATE function becomes simpler and tends
Figure 5: Performance of H-learner across different $\lambda$ values for Setup A.
to favor direct regularization. We observe that the H-learner adapts to this inductive bias by yielding a larger optimal $\lambda$ as the number of shared features increases. Similar performance patterns are observed in Setups B and C, and results are provided in the supplement.
@@ 5.3 Benchmark Evaluation
Here we present results on IHDP and ACIC 2016 benchmark datasets.
Baselines. We consider T-learner and TARNet as baselines for indirect learners, and X-learner, DR-learner, and IPW-learner as baselines for direct learners. Additionally, we include three state-of-the-art regularization strategies for indirect learners: (1) Weight Regularization: TARNet-WR, a variant of TARNet with an explicit regularization that penalizes differences between the weights of the two output heads. (2) Reparametrization: OffsetNet, a neural network–based indirect learner that achieves implicit regularization through the reparameterization $\mu_{1}(x)=\mu_{0}(x)+\tau(x)$, enabling direct estimation of $\tau(x)$ as an offset from $\mu_{0}(x)$ [10, 20]. (3) Structural Regularization: FlexTENet, a neural network architecture that imposes regularization through flexible feature sharing across layers [10].
Comparison to other regularization strategies: Note that all other existing regularization strategies fail to perform well on these benchmarks. As discussed in Section 4.2, these strategies rely on the assumption that reparameterization or sharing of weights simplifies the learning problem—an assumption that only holds when the underlying treatment effect is small. Since the assumption fails to hold in practical scenarios, using these strategies can actually harm performance. For example, TARNet with weight regularization performs worse than the standard TARNet model. This further highlights the robustness of H-learner in handling diverse data conditions without relying on structural assumptions.
Performance breakdown by DGP characteristics: We further break down the performance on ACIC 2016 by specific DGP characteristics, with detailed results provided in the supplement. The H-learner consistently achieves the best performance across diverse settings, including variations in response models, treatment effect heterogeneity, confounding strength, and covariate overlap.
Table 2: Within-sample and out-of-sample results on benchmark datasets IHDP and ACIC 2016.
@@
TARNet $0.81\pm.02$ $0.93\pm.03$ 1.66±.03plus-or-minus1.66.031.$6$03 1.77±.04plus-or-minus1.77.041.$7$04
@@
TARNet-WR $1.19\pm.03$ 1.33±.04plus-or-minus1.33.041.$3$04 $1.89\pm.04$ $1.97\pm.04$
OffsetNet 2.11±.08plus-or-minus2.11.082.$1$08 $2.18\pm.09$ $2.10\pm.05$ $2.15\pm.05$
@@
H-learner (X) $\bm{0.78\pm.01}$ $\bm{0.88\pm.02}$ $\bm{1.56\pm.03}$ $\bm{1.67\pm.04}$

### v2

@@ 4 Related Work
4.1 Inconsistent Performance of Indirect and Direct Learners
The divergence in performance between indirect and direct learners has been consistently observed across settings in the literature. In particular, prior work has shown that no single meta-learner consistently outperforms others across all settings (Künzel et al., 2019). Both theoretical and empirical studies show that their relative performance varies substantially with the sparsity and smoothness of the response surfaces (Curth and Van der Schaar, 2021). Acharki et al. (2023) extended the comparison of meta-learners to the non-binary treatment setting, theoretically highlighting the bias–variance trade-off in the multi-treatment setting. These inconsistencies highlight the importance of developing a more robust approach, as the true DGP is often unknown in practice, and relying on either method can result in unreliable performance.
4.2 Regularization Strategies for CATE Estimations
Several other regularization strategies have been proposed to mitigate “regularization-induced confounding”. Weight regularization penalizes discrepancies between the weights of the outcome heads to encourage parameter sharing (Hahn et al., 2018). Reparameterization implicitly introduces regularization by modeling $\mu_{1}(x)=\mu_{0}(x)+\tau(x)$, allowing the treatment effect $\tau(x)$ to be predicted directly by one head as an offset from $\mu_{0}(x)$ (Imai and Ratkovic, 2013; Curth and van der Schaar, 2021). Structural regularization was proposed with FlexTENet, a specialized neural architecture that enforces regularization through flexible feature sharing across layers (Curth and van der Schaar, 2021).
@@ 4.2 Regularization Strategies for CATE Estimations
5 Theoretical Analysis
In this section, we analyze the statistical properties of the H-learner. We focus on the linear setup, where the H-learner admits a closed-form solution. We then characterize conditions under which H-learner achieves strictly lower prediction risk than either estimator alone. Extending to nonlinear function classes does not yield a tractable closed form, so we defer to experiments in Section 6 for empirical validation. All proofs are deferred to Appendix B.
5.1 Closed-Form Solution of the H-Learner in Linear Models
Throughout this section, we use $X\in\mathbb{R}^{n\times d}$ to denote the covariate matrix and $X_{1}\in\mathbb{R}^{n_{1}\times d}$ and $X_{0}\in\mathbb{R}^{n_{0}\times d}$ denote the submatrices of $X$ corresponding to treated and control groups. Let $\tau(x)$ denote the true CATE, and define $\theta^{\star}=\arg\min_{\theta}\,\mathbb{E}\big[(x^{\top}\theta-\tau(x))^{2}\big]$ as the true treatment effect parameter in the linear $\tau(x)$ case, or the best linear projection of $\tau(x)$ onto the span of $x$ in the nonlinear case over the feature space. We denote by $\hat{\theta}_{\mathrm{ind}}$, $\hat{\theta}_{\mathrm{dir}}$, and $\hat{\theta}_{H}$ the OLS parameter estimates produced by the indirect learner, direct learner, and H-learner, respectively. For ease of notation, let $G_{1}=X_{1}^{\top}X_{1},\ G_{0}=X_{0}^{\top}X_{0},\ G=X^{\top}X$.
Theorem 5.1 (Exact H-learner). Under the linear setup, the H-learner OLS estimator admits a closed-form solution given by a matrix-weighted combination of the indirect and direct OLS estimators.
$\hat{\theta}_{H}=(I-W)\hat{\theta}_{\mathrm{ind}}+W\hat{\theta}_{\mathrm{dir}},$
where $W=\lambda A\big[(1-\lambda)I+\lambda A\big]^{-1},\ A:=(G_{1}^{-1}+G_{0}^{-1})\,G.$
Theorem 5.1 shows that the H-learner forms a matrix-weighted linear combination with a weight that depends on the regularization parameter $\lambda$ and the geometry encoded in the matrix $A$. When $\lambda=0$, the H-learner reduces to the indirect learner, and when $\lambda=1$, it equals the direct learner. In general, this weighting operates in a direction-wise manner across the feature space. Let $\{\mu_{i}\}_{i=1}^{d}$ be the eigenvalues of $A$. Then the eigenvalues of $W$ are $\left\{\frac{\lambda\mu_{i}}{(1-\lambda)+\lambda\mu_{i}}\right\}_{i=1}^{d}$, which are strictly increasing in both $\lambda$ and $\mu_{i}$. This highlights the adaptivity of the H-learner: Under poor overlap or in unbalanced settings with a small treatment arm, the eigenvalues of $A$ tend to increase, as the Gram matrix $G_{1}$ or $G_{0}$ corresponding to the smaller treatment group becomes ill-conditioned. Consequently, the eigenvalues of $W$ increase, shifting more weight toward the direct approach, where indirect learners are more susceptible to bias induced by separate regularization.
We next analyze how this interpolation translates into improved prediction risk through a bias–variance decomposition.
5.2 Bias–Variance Analysis of the H-Learner
Let $e_{\mathrm{ind}}=\hat{\theta}_{\mathrm{ind}}-\theta^{\star}$ and $e_{\mathrm{dir}}=\hat{\theta}_{\mathrm{dir}}-\theta^{\star}$ denote the estimation errors of the indirect and direct learners, respectively. We then define their biases $b_{\mathrm{ind}}=\mathbb{E}[e_{\mathrm{ind}}]$ and $b_{\mathrm{dir}}=\mathbb{E}[e_{\mathrm{dir}}]$, and their covariance matrices $\Sigma_{\mathrm{ind}}=\operatorname{Var}(e_{\mathrm{ind}})$, $\Sigma_{\mathrm{dir}}=\operatorname{Var}(e_{\mathrm{dir}})$, $\Sigma_{\mathrm{ind,dir}}=\operatorname{Cov}(e_{\mathrm{ind}},e_{\mathrm{dir}})$. The mean squared error (MSE) can be expressed as $\mathrm{MSE}_{\mathrm{ind}}=\mathbb{E}[\|e_{\mathrm{ind}}\|_{2}^{2}]$ and $\mathrm{MSE}_{\mathrm{dir}}=\mathbb{E}[\|e_{\mathrm{dir}}\|_{2}^{2}]$.
To facilitate the analysis, we start by focusing on the scalar path $W=\omega I$, which leads to closed-form conditions under which the H-learner outperforms both endpoints. Since $\min_{W}\mathrm{MSE}_{H}(W)\leq\min_{\omega\in[0,1]}\mathrm{MSE}_{H}(\omega I)$, any improvement achieved along this scalar path provides a sufficient condition, as the full matrix optimization yields no higher risk and thus will also improve upon both endpoints.
Assumption 5.2. We assume that the indirect and direct learners are trained via cross-fitting (e.g., on independent folds), so that $\Sigma_{\mathrm{ind,dir}}=0.$
Theorem 5.3 (MSE expansion and optimal weight). Under Assumptions 5.2, the MSE of the H-learner estimator is
$\mathrm{MSE}_{\mathrm{H}}=\mathbb{E}\big[\lVert\hat{\theta}_{H}-\theta^{\star}\rVert_{2}^{2}\big]=(1-\omega)^{2}\,\mathrm{MSE}_{\mathrm{ind}}\\ +\,\omega^{2}\,\mathrm{MSE}_{\mathrm{dir}}+\,2\omega(1-\omega)\,D,$
where $D=b_{\mathrm{ind}}^{\top}b_{\mathrm{dir}}$.
The optimal convex weight that minimizes $\mathrm{MSE}_{\mathrm{H}}$ is
$\omega^{*}=\min\!\left\{\,1,\;\max\!\left\{\,0,\;\frac{\mathrm{MSE}_{\mathrm{ind}}-D}{\mathrm{MSE}_{\mathrm{ind}}+\mathrm{MSE}_{\mathrm{dir}}-2D}\,\right\}\right\}.$
Theorem 5.3 decomposes the H-learner’s MSE into a weighted sum of the direct and indirect errors together, and characterizes the optimal weight $\omega^{*}$. Consequently, the H-learner strictly outperforms both learners whenever $0<\omega^{*}<1$. We next present two corollaries that provide sufficient conditions for this to hold.
Corollary 5.4. If $b_{\mathrm{ind}}^{\top}b_{\mathrm{dir}}<0$, then $\mathrm{MSE}_{\mathrm{H}}<\min\!\left\{\mathrm{MSE}_{\mathrm{ind}},\;\mathrm{MSE}_{\mathrm{dir}}\right\}$.
Corollary 5.5. If $\|b_{\mathrm{ind}}\|_{2}\geq\|b_{\mathrm{dir}}\|_{2}$ and $\operatorname{tr}(\Sigma_{\mathrm{dir}})>b_{\mathrm{dir}}^{\top}(b_{\mathrm{ind}}-b_{\mathrm{dir}})$, then $\mathrm{MSE}_{\mathrm{H}}<\min\!\left\{\mathrm{MSE}_{\mathrm{ind}},\;\mathrm{MSE}_{\mathrm{dir}}\right\}$.
Corollary 5.4 shows that when the biases point in opposite directions, the H-learner benefits from their cancellation. When they point in the same direction, Corollary 5.5 makes explicit the bias–variance trade-off between indirect and direct learners. It implies that whenever the indirect learner has a larger bias and the direct learner’s variance is sufficiently large relative to this bias gap, then the H-learner strictly outperforms both individual estimators.
We further illustrate this with a 1D example in Figure 3. The plot shows the MSE of the H-learner as a function of the convex weight $\omega$, together with its decomposition into bias and variance. As expected, the indirect learner exhibits higher bias, while the direct learner suffers from higher variance. H-learner achieves strictly lower MSE than either baseline by reducing variance below both while maintaining a small level of bias.
Figure 3: MSE and bias-variance decomposition of H-learner across convex weight $\omega$. Empirical optimum $\omega^{\star}$ aligns with the theoretical optimal point. H-learner with $\omega^{\star}$ achieves lower MSE than either baseline.
@@ 6 Experiments
In this section, we evaluate the performance of H-learner through semi-synthetic experiments and established benchmarks. Section 6.1 outline the datasets and experimental setup. Section 6.2 highlight the inconsistent performance between indirect and direct learners using custom-designed semi-synthetic setups and demonstrate the robustness of H-learner. In Section 6.3, we present results on widely used benchmark datasets, further highlighting the improved performance of H-learner over existing baselines.
Figure 4: PEHE results for three semi-synthetic scenarios. In all setups, we observe performance trade-offs between TARNet (indirect learner) and X-learner (direct learner), while the proposed H-learner consistently outperforms both counterparts.
@@ 6.1 Setup
Datasets. We evaluate performance using two widely adopted benchmark datasets: IHDP and ACIC 2016, both commonly used in the literature on heterogeneous treatment effect estimation (Shalit et al., 2017; Alaa and Van Der Schaar, 2017; Shi et al., 2019; Lee et al., 2020; Curth and Van der Schaar, 2021). Following prior work, we use a 63/27/10 train/validation/test split for both datasets.
IHDP. The IHDP dataset is based on covariates from the Infant Health and Development Program, a randomized study assessing the effect of specialist home visits on cognitive outcomes for premature infants (Hill, 2011). Confounding and treatment imbalance are introduced by removing a subset of treated units. The resulting dataset includes 747 observations with 25 covariates. Following prior work, we evaluate performance by averaging over 1,000 realizations from setting “A” implemented in the NPCI package (Dorie, 2016).
ACIC 2016. The ACIC 2016 datasets are derived from the Collaborative Perinatal Project and were introduced as part of the Atlantic Causal Inference Competition (Dorie et al., 2019). This benchmark provides a comprehensive evaluation that varies in functional complexity, confounding, overlap, and treatment effect heterogeneity. It includes 77 distinct DGPs, with 4,802 observations and 79 covariates. Following (Lee et al., 2020), we remove rows where the POs below the 1st and above the 99th percentile to eliminate outliers and evaluate performance by averaging over 5 random realizations per setting.
Evaluations. We evaluate performance using Precision in Estimation of Heterogeneous Effect (PEHE) (Hill, 2011). The PEHE loss is defined as $\epsilon_{\text{PEHE}}=\frac{1}{n}\mbox{$\sum_{i=1}^{n}$}(\widehat{\tau}(x_{i})-\tau(x_{i}))^{2}$, which corresponds to the mean squared error in estimating the CATE. We report performance on both in-sample (training and validation) and out-of-sample (test) data. Note that in-sample evaluation is non-trivial, as only the factual outcomes are observed during training, while the true CATE $\tau(x)$ is never observed.
@@ 6.2 Semi-Synthetic Experiments
Experiment setup. In all setups, we use covariates from the IHDP dataset (n=747,d=25n=747,\ d=25) but design custom response surfaces as follows.
Let $S_{0}$ and $S_{1}$ denote sets of randomly selected features used to construct $\mu_{0}(x)$ and $\mu_{1}(x)$, respectively, and let $S=S_{0}\cup S_{1}$. We then simulate the response surfaces and the treatment assignment as follows:
[…]
$\mathbb{P}(T_{i}=1\mid X_{i})=\sigma\!\left(\alpha\cdot\mbox{$\sum_{\begin{subarray}{c}j\in S\end{subarray}}$}\ \beta_{j}X_{j}\right),$
@@
In Setup A, we vary the relative complexity of the CATE and PO functions by changing the proportion of shared features between $S_{0}$ and $S_{1}$ in $\{10\%,30\%,50\%,70\%,90\%\}$, where a greater number of shared features implies a simpler CATE function. In Setup B, we vary group imbalance by adjusting the proportion of treated units in the dataset to $\{20\%,30\%,40\%,50\%\}$, where a higher percentage indicates more balanced treatment and control groups. In Setup C, we introduce varying levels of confounding in the treatment assignment function by adjusting $\alpha\in\{0,0.2,0.4,0.6,0.8\}$, where larger values of $\alpha$ correspond to stronger confounding. All simulations are averaged over 80 runs, with features in $S_{0}$ and $S_{1}$ randomly selected in each run.
Results. Figure 4 presents results from the three simulated scenarios. We show the performance of the H-learner using X-learner pseudo-outcomes in its first stage and compare it against its constituent components: the X-learner (direct learner) and TARNet (indirect learner). Results using alternative learners are provided in Appendix C.
Setup A shows the performance trade-off through the relative complexity of the POs and the CATE. As the proportion of shared features increases, the PO functions become more similar, which leads to a simpler CATE surface. Correspondingly, we observe that the X-learner only outperforms TARNet when the proportion of shared features is high, as its constructed pseudo-outcomes target a simpler function, but underperforms when the shared feature proportion is low.
Figure 5: Performance of H-learner across different $\lambda$ values for Setup A. Optimal performance is consistently achieved at intermediate $\lambda$ values.
Setup B demonstrates the performance trade-off driven by the treatment–control imbalance. As the proportion of treated units decreases, the imbalance in sample sizes between the treatment and control groups becomes more severe. This imbalance disproportionately affects indirect learners, as their separate outcome models are trained on increasingly unbalanced samples and are more susceptible to “regularization-induced confounding”. Empirically, TARNet performs better under balanced treatment but underperforms when only 20% of units receive treatment.
@@
In contrast, across all varying setups, the H-learner consistently outperforms both its direct and indirect counterparts by adjusting its inductive bias to better align with the underlying data, demonstrating robustness by unifying the strengths of both learners.
Effect of regularization parameter $\lambda$. We further investigate the performance of H-learner across different $\lambda$ values. Figure 5 shows how the test PEHE varies with $\lambda$ in Setup A, across different levels of feature sharing: low (10%), medium (50%), and high (90%).
[…]
@@ 6.3 Benchmark Evaluation
Here we present results on the IHDP and ACIC 2016 benchmark datasets, comparing the H-learner with baseline meta-learners and regularization strategies from the literature.
Table 2: Within-sample and out-of-sample PEHE results on benchmark datasets IHDP and ACIC 2016. Entries marked with indicate statistically significant improvements over all baselines based on one-sided paired t-tests.
@@
In-sample $\sqrt{\epsilon_{\text{PEHE}}}$ Out-of-sample $\sqrt{\epsilon_{\text{PEHE}}}$ In-sample $\sqrt{\epsilon_{\text{PEHE}}}$ Out-of-sample $\sqrt{\epsilon_{\text{PEHE}}}$
@@
TARNet $0.81\pm.02$ $0.93\pm.03$ $1.66\pm.03$ $1.77\pm.04$
@@
TARNet-WR $1.19\pm.03$ $1.33\pm.04$ $1.89\pm.04$ $1.97\pm.04$
OffsetNet $2.11\pm.08$ $2.18\pm.09$ $2.10\pm.05$ $2.15\pm.05$
@@
H-learner (X) $\bm{0.78\pm.01}^{\star}$ $\bm{0.88\pm.02}^{\star}$ $\bm{1.56\pm.03}^{\star}$ $\bm{1.67\pm.04}^{\star}$
Baselines. We consider T-learner and TARNet as baselines for indirect learners, and X-learner, DR-learner, and IPW-learner as baselines for direct learners. Additionally, we include three state-of-the-art regularization strategies: (1) Weight Regularization: TARNet-WR, a variant of TARNet with an explicit regularization that penalizes differences between the weights of the two output heads (Hahn et al., 2018; Curth and van der Schaar, 2021). (2) Reparametrization: OffsetNet, a neural network–based indirect learner that achieves implicit regularization through the reparameterization $\mu_{1}(x)=\mu_{0}(x)+\tau(x)$, enabling direct estimation of $\tau(x)$ as an offset from $\mu_{0}(x)$ (Imai and Ratkovic, 2013; Curth and van der Schaar, 2021). (3) Structural Regularization: FlexTENet, a neural network architecture that imposes regularization through flexible feature sharing across layers (Curth and van der Schaar, 2021).
Figure 6: Radar chart of ACIC performance by DGP characteristics. PEHE is min–max normalized so that higher values indicate better performance, with 1 denoting the best. Only the top method per class is shown. H-learner performs best across all settings.
Comparison to other regularization strategies: Note that all other existing regularization strategies fail to perform well on these benchmarks. As discussed in Section 4.2, these strategies rely on the assumption that reparameterization or sharing of weights simplifies the learning problem—an assumption that only holds when the treatment effect is small. Since the assumption fails to hold in practical scenarios, using these strategies can actually harm performance. For example, TARNet with weight regularization performs worse than the standard TARNet model. This further highlights the robustness of H-learner in handling diverse data conditions without relying on any structural assumptions.
Performance breakdown by DGP characteristics: We further analyze performance on ACIC 2016 by decomposing results along specific DGP characteristics, as summarized in Figure 6. The H-learner consistently achieves the best performance across a wide range of settings, including variations in response models, treatment effect heterogeneity, confounding strength, and covariate overlap. Complete results are provided in Appendix D.